# Tight Guarantees for Cut-Relative Survivable Network Design via a Decomposition Technique

### v2

@@ 1 Introduction
$\mathsf{SNDP}$ imposes an absolute level of fault-tolerance in a solution subgraph $H$, by ensuring that for any bounded-size fault-set $F\subseteq E$ of edges that may fail, the graph $H-F$ retains some connectivity properties. More precisely, due to Menger’s theorem, or the max-flow min-cut theorem, the feasibility condition on $H$ can be stated equivalently as follows. For every $i\in[k]$, and every fault-set $F\subseteq E$ with $|F|<r_{i}$:
@@
    (Path-version) there is an $s_{i}$-$t_{i}$ path in $H-F$;
@@
Recently, motivated by work on fault-tolerance in graph spanners, Dinitz, Koranteng, and Kortsartz [8], proposed a variant of $\mathsf{SNDP}$ that aims to impose a relative level of fault-tolerance with respect to the graph $G$. The idea is that even if the $\mathsf{SNDP}$ instance in infeasible, because $G$ itself does not possess the required level of fault-tolerance, one should not have to completely abandon the goal of fault-tolerance: it is still meaningful and reasonable to seek a solution that is “as fault-tolerant as $G$”, and in this sense is fault-tolerant relative to $G$.
@@
for every fault-set $F$ (valid for $\mathsf{SNDP}$),  $H-F$ and $G-F$ have similar connectivity. (*)
To elaborate, in $\mathsf{SNDP}$, if there is even one fault-set $F^{\prime}$ under which $G-F^{\prime}$ fails to have the required connectivity, i.e., $|F^{\prime}|<r_{i}$ and $F^{\prime}\supseteq\delta_{G}(S)$ for some $s_{i}$-$t_{i}$ cut $S$, then “all bets are off”; we declare that the instance in infeasible. But declaring infeasibility here feels unsatisfactory because we allow one specific “problematic” fault-set $F^{\prime}$ to block us from obtaining any kind of fault-tolerance. In contrast, (*1) aims to provide a per-fault-set guarantee, by asking for a solution $H$ that functions as well as $G$ in terms of connectivity under the failure of any fault-set $F$ (considered for $\mathsf{SNDP}$). We can formalize “$H-F$ and $G-F$ have similar connectivity” in two ways, via paths or cuts, and this gives rise to the following two problem definitions.
@@
    Path-relative $\mathsf{SNDP}$: $H\subseteq E$ is a feasible solution if for every $i\in[k]$ and $F\subseteq E$ with $|F|<r_{i}$,   $G-F$ has an $s_{i}$-$t_{i}$ path   $\implies$   $H-F$ has an $s_{i}$-$t_{i}$ path.
@@
Perhaps surprisingly, path-relative $\mathsf{SNDP}$ and cut-relative $\mathsf{SNDP}$ are not equivalent, even when the $\mathsf{SNDP}$ instance involves a single $s$-$t$ pair. It is not hard to show that if $H$ is feasible for $\mathsf{CR}\text{-}\mathsf{SNDP}$, it is also feasible for path-relative SNDP [^1] Fix any $i\in[k]$ and $F\subseteq E$ with $|F|<r_{i}$. If $G-F$ has an $s_{i}$-$t_{i}$ path, then $\delta_{G-F}(S)\neq\emptyset$ for every $s_{i}$-$t_{i}$ cut. So since $H$ is feasible for cut-relative $\mathsf{SNDP}$, we have $\delta_{H-F}(S)\neq\emptyset$ for every $s_{i}$-$t_{i}$ cut, which implies that $H-F$ has an $s_{i}$-$t_{i}$ path. but the converse fails to hold; see Appendix A for an example.
@@
We arrived at $\mathsf{CR}\text{-}{f}\text{-}\mathsf{NDP}$, which can be seen as a fault-tolerant cut-covering problem, as a technically natural extension of $k\text{-}\mathsf{EFTS}$. We show below that $\mathsf{CR}\text{-}{f}\text{-}\mathsf{NDP}$ offers a surprising amount of modeling power, and, in particular, allows one to capture stronger forms of relative fault-tolerance compared to $k\text{-}\mathsf{EFTS}$. In $k\text{-}\mathsf{EFTS}$, there is a sharp relative-fault-tolerance threshold at $k$: if $H$ is feasible, then $G-F$ and $H-F$ have the same components whenever $|F|<k$, but there are no guarantees for larger fault-sets. With $\mathsf{CR}\text{-}{f}\text{-}\mathsf{NDP}$, one can capture a weaker guarantee also for other fault-sets, which allows for a more graceful degradation of relative fault-tolerance as $|F|$ increases.
@@
One way of modeling graceful relative-fault-tolerance degradation is as follows: given a non-increasing function $\tau:\mathbb{Z}_{+}\mapsto\mathbb{Z}_{+}$, we seek a min-cost subgraph $H$ such that for every fault-set $F$, $G-F$ and $H-F$ have exactly the same components with at most $\tau(|F|)$ nodes. If $G$ models a communication network where connected nodes can communicate with each other, then this yields the desirable guarantee that, post-faults, if a node can communicate with more than $\tau(|F|)$ nodes in $G$ then the same holds for $H$; if this number is at most $\tau(|F|)$, then post-faults, it can communicate with the same nodes in $G$ and $H$. Note that if $\tau(k-1)\geq n=|V|$, then $H$ is feasible for $k\text{-}\mathsf{EFTS}$; by adding more “levels” to $\tau$, we can obtain fault-tolerance guarantees for larger fault-sets.
@@
Generalizing Example 1.2 substantially, suppose along with $\tau$, we have a monotone function , i.e., if $T\subseteq S$. We now seek a subgraph $H$ such that for every fault-set $F$ and $S\subseteq V$ with , we have that $S$ is (the node-set of) a component of $G-F$ iff it is a component of $H-F$. Similar to Example 1.2, we can model this as $\mathsf{CR}\text{-}{f}\text{-}\mathsf{NDP}$ by defining $f(S):=\min\,\{\ell:\tau(\ell)<$, for $S\neq\emptyset$ (Theorem 3.3).
This setup yields fault-tolerance degradation depending on monotone properties of components other than just their size, which creates a rich space of problems. For example, suppose is the weak-diameter of $S$. Setting $\tau(\ell)=n$ if $\ell<k$, and some $t<n$ for larger $\ell$, we seek a $k\text{-}\mathsf{EFTS}$ solution $H$ such that for $F\subseteq E$ with $|F|\geq k$, $G-F$ and $H-F$ have the same components of weak-diameter at most $t$.
@@ 1.1 Our contributions
We introduce a little notation to describe this. Let $\mathsf{CR}\text{-}{(h,D)}\text{-}\mathsf{NDP}$ denote $\mathsf{CR}\text{-}{f}\text{-}\mathsf{NDP}$ on graph $D$ with base cut-requirement function $h$. For $S\subseteq V$, define its deficiency to be $f(S)-|\delta_{G}(S)|$; we say that $S$ is a small cut if its deficiency is positive (i.e., $|\delta_{G}(S)|<f(S)$). We use $G[{S}]$ to denote the subgraph induced by $S$. Let $\overline{S}$ denote $V-S$.
@@
Iterative rounding and relaxation derives its power from the fact that an extreme-point solution to an LP-relaxation of the problem can be defined using a structured family of tight constraints, such as, often a laminar family of cut constraints in the case of network-design problems. As noted earlier, this key property does not hold for cut-relative $\mathsf{SNDP}$ (and path-relative $\mathsf{SNDP}$). Recently, various works have considered network-design problems that give rise to cut constraints that are less structured and not quite uncrossable; see [Bansal25, 3, 6] and the references therein. Our work can be seen as furthering this research direction.
@@ 2 Preliminaries and notation
Recall that we are given an undirected graph $G=(V,E)$ with nonnegative edge costs $\{c_{e}\}_{e\in E}$. For a subset $S\subseteq V$ and subset $Z\subseteq E$, which we will interchangeably view as the subgraph $(V,Z)$ of $G$, we use $\delta_{Z}(S)$ to denote $\delta_{G}(S)\cap Z$, i.e., the edges of $Z$ on the boundary of $S$. Recall that $\overline{S}$ denotes $V-S$, and $G[{S}]$ denotes the subgraph induced by $S$. A cut-requirement function on $G$ is a function $f:2^{V}\mapsto\mathbb{Z}$. (We allow $f(S)$ to be negative chiefly for notational simplicity; this does not affect anything, as we will only examine sets for which $f(S)\geq 0$.) Recall the following network-design problems.
@@ 4. $\bullet$
Let $f:2^{V}\mapsto\mathbb{R}$. We say that $f$ is weakly supermodular if for any two node-sets $A,B\subseteq V$, we have $f(A)+f(B)\leq\max\bigl{\{}f(A\cap B)+f(A\cup B),\,f(A-B)+f(B-A)\bigr{\}}$. We say that $f$ is symmetric if $f(S)=f(V-S)$ for all $S\subseteq V$, and is normalized if $f(\emptyset)=f(V)=0$.
@@
    Let $X\subseteq V$, and $\mathcal{C}$ be a collection of subsets of $V-X$ closed under taking set intersections, unions, differences, and complements with respect to $V-X$, i.e., if $S\in\mathcal{C}$ then $V-X-S\in\mathcal{C}$, and $S,T\in\mathcal{C}$ implies that $\{S\cap T,S\cup T,S-T,T-S\}\subseteq\mathcal{C}$.
@@ 1. $\bullet$
    $X=A$,  $Y=V-B$. We have $X\cap Y=A-B$, and $X\cup Y=V-(B-A)$, so $f^{\mathsf{sym}}(X\cup Y)=f^{\mathsf{sym}}(B-A)$. Also, $X-Y=A\cap B$, and $Y-X=V-(A\cup B)$, so $f^{\mathsf{sym}}(Y-X)=f^{\mathsf{sym}}(A\cup B)$. So (1) holds.
@@ 2. $\bullet$
    $X=V-A$,  $Y=B$. This is completely symmetric to the above case. We have $f^{\mathsf{sym}}(X\cap Y)=f^{\mathsf{sym}}(B-A)$ and $f^{\mathsf{sym}}(X\cup Y)=f^{\mathsf{sym}}(A-B)$. Also, $f^{\mathsf{sym}}(X-Y)=f^{\mathsf{sym}}(A\cup B)$ and $f^{\mathsf{sym}}(Y-X)=f^{\mathsf{sym}}(A\cap B)$.
@@ 3. $\bullet$
    $X=V-A$,  $Y=V-B$. We have $X\cap Y=V-(A\cup B)$, so $f^{\mathsf{sym}}(X\cap Y)=f^{\mathsf{sym}}(A\cup B)$, and $X\cup Y=V-(A\cap B)$, so $f^{\mathsf{sym}}(X\cup Y)=f^{\mathsf{sym}}(A\cap B)$. Also, $X-Y=B-A$ and $Y-X=A-B$.
@@
Now we show that $g_{2}$ is weakly supermodular. Consider $A,B\subseteq X$. If any of the sets $A-B,B-A,A\cap B,X-(A\cup B)$ are empty, then $g_{2}(A)+g_{2}(B)=g_{2}(A\cap B)+g_{2}(A\cup B)$ or $g_{2}(A)+g_{2}(B)=g_{2}(A-B)+g_{2}(B-A)$. The former holds when $A-B$ or $B-A$ is empty; the latter holds when $A\cap B$ or $X-(A\cup B)$ is empty, where we also utilize the fact that $g_{2}$ is symmetric. So assume otherwise. Then, for every $S\in\{A,B,A\cap B,A\cup B,A-B,B-A\}$, we have $g_{2}(S)=\max_{T\in\mathcal{C}}f(S\cup T)$.
@@ 2.1 LP-relaxation for $\mathsf{CR}\text{-}{f}\text{-}\mathsf{NDP}$
We show that for the $\mathsf{CR}\text{-}\mathsf{SNDP}$ instance shown in Fig. 1, the extreme-point solution $\widehat{x}$ shown in the figure, cannot be expressed as the solution to any laminar family of tight cut constraints (2) (along with tight edge constraints (3)). The $\mathsf{SNDP}$ instance is $s$-$t$ $2$-edge connectivity: that is, the base cut-requirement function is given by $f(S)=2$ if $S$ is an $s$-$t$ cut, and $f(S)=0$ otherwise.
@@ 3 Modeling power of $\mathsf{CR}\text{-}{f}\text{-}\mathsf{NDP}$
    $H$ is feasible for $\mathsf{CR}\text{-}{f^{\mathsf{grace}}}\text{-}\mathsf{NDP}$ iff for every fault-set $F\subseteq E$, and every $S\subseteq V$ with $|S|\leq\tau(|F|)$, $S$ is (the node-set of) a component of $G-F$ iff it is a component of $H-F$.
@@
To prove the feasibility characterization in part (b), we note that the function $f^{\mathsf{grace}}$ is constructed so that for any $S\subseteq V$ and $F\subseteq E$, we have $|S|\leq\tau(|F|)$ iff $f^{\mathsf{grace}}(S)>|F|$, so that $|\delta_{H}(S)|\geq f^{\mathsf{grace}}(S)$ implies that $S$ is not a component of $H-F$. Suppose $H$ is such that $|\delta_{H}(S)|\geq\min\bigl{\{}f^{\mathsf{grace}}(S),|\delta_{G}(S)|\bigr{\}}$ for all $S\subseteq V$. Consider any fault-set $F\subseteq E$ and any $S\subseteq V$ such that $|S|\leq\tau(|F|)$. By the above, if $S$ is a component of $H-F$, then $|\delta_{H}(S)|<f^{\mathsf{grace}}(S)$, and so $S$ must also be a component of $G-F$. But this also implies that if $S$ is a component of $G-F$, then it is a component of $H-F$; otherwise, there is some $A\subsetneq S$ that is a component of $H-F$ with $|A|<|S|\leq\tau(|F|)$, and so $A$ must be a component of $G-F$, which is not the case.
Conversely, suppose that for every $F\subseteq E$, and $S\subseteq V$ with $|S|\leq\tau(|F|)$, we have that $S$ is a component of $G-F$ iff it is a component of $H-F$. If $|\delta_{H}(S)|<\min\bigl{\{}f^{\mathsf{grace}}(S),|\delta_{G}(S)|\bigr{\}}$ for some $S\subseteq V$, then take $F=\delta_{H}(S)$. Since $f^{\mathsf{grace}}(S)>|F|$, we have $|S|\leq\tau(|F|)$. But $S$ is a component of $H-F$ and not $G-F$, which yields a contradiction.
@@
    $H$ is feasible for $\mathsf{CR}\text{-}{f}\text{-}\mathsf{NDP}$ iff for every fault-set $F\subseteq E$, and every $S\subseteq V$ with , $S$ is (the node-set of) a component of $G-F$ iff it is a component of $H-F$.
@@
For part (b), similar to before, we have that iff $f(S)>|F|$. Suppose $H$ is feasible for $\mathsf{CR}\text{-}{f}\text{-}\mathsf{NDP}$. Consider any fault-set $F\subseteq E$ and any $S\subseteq V$ be such that . By the above, $f(S)>|F|$, so if $S$ is a component of $H-F$, then it must also be a component of $G-F$. Again, this also implies that if $S$ is a component of $G-F$, then it is a component of $H-F$: otherwise, there is some $A\subsetneq S$ that is a component of $H-F$, and , which means that $A$ is a component of $G-F$.
Conversely, suppose that for every $F\subseteq E$, and $S\subseteq V$ with , we have that $S$ is a component of $G-F$ iff it is a component of $H-F$. If $|\delta_{H}(S)|<\min\bigl{\{}f(S),|\delta_{G}(S)|\bigr{\}}$ for some $S\subseteq V$, then take $F=\delta_{H}(S)$. Since $f(S)>|F|$, we have . But $S$ is a component of $H-F$ and not $G-F$, which yields a contradiction.

### v3

@@ 1 Introduction
$\mathsf{SNDP}$ imposes an absolute level of fault-tolerance in a solution subgraph $H$, by ensuring that for any bounded-size fault-set $F\subseteq E$ of edges that may fail, the graph HFH-F retains some connectivity properties. More precisely, due to Menger’s theorem, or the max-flow min-cut theorem, the feasibility condition on $H$ can be stated equivalently as follows. For every $i\in[k]$, and every fault-set $F\subseteq E$ with $|F|<r_{i}$:
@@
    (Path-version) there is an $s_{i}$-$t_{i}$ path in HFH-F;
@@
Recently, motivated by work on fault-tolerance in graph spanners, Dinitz, Koranteng, and Kortsartz [8], proposed a variant of $\mathsf{SNDP}$ that aims to impose a relative level of fault-tolerance with respect to the graph $G$. The idea is that even if the $\mathsf{SNDP}$ instance in infeasible, because $G$ itself does not possess the required level of fault-tolerance, one should not have to completely abandon the goal of fault-tolerance: it is still meaningful and reasonable to seek a solution that is “as fault-tolerant as GG”, and in this sense is fault-tolerant relative to $G$.
@@
for every fault-set $F$ (valid for $\mathsf{SNDP}$),  HFH-F and GFG-F have similar connectivity. (*)
To elaborate, in $\mathsf{SNDP}$, if there is even one fault-set $F^{\prime}$ under which $G-F^{\prime}$ fails to have the required connectivity, i.e., $|F^{\prime}|<r_{i}$ and $F^{\prime}\supseteq\delta_{G}(S)$ for some $s_{i}$-$t_{i}$ cut $S$, then “all bets are off”; we declare that the instance in infeasible. But declaring infeasibility here feels unsatisfactory because we allow one specific “problematic” fault-set $F^{\prime}$ to block us from obtaining any kind of fault-tolerance. In contrast, (*1) aims to provide a per-fault-set guarantee, by asking for a solution $H$ that functions as well as $G$ in terms of connectivity under the failure of any fault-set $F$ (considered for $\mathsf{SNDP}$). We can formalize “HFH-F and GFG-F have similar connectivity” in two ways, via paths or cuts, and this gives rise to the following two problem definitions.
@@
    Path-relative $\mathsf{SNDP}$: $H\subseteq E$ is a feasible solution if for every $i\in[k]$ and $F\subseteq E$ with $|F|<r_{i}$,   GFG-F has an $s_{i}$-$t_{i}$ path   $\implies$   HFH-F has an $s_{i}$-$t_{i}$ path.
@@
Perhaps surprisingly, path-relative $\mathsf{SNDP}$ and cut-relative $\mathsf{SNDP}$ are not equivalent, even when the $\mathsf{SNDP}$ instance involves a single $s$-$t$ pair. It is not hard to show that if $H$ is feasible for $\mathsf{CR}\text{-}\mathsf{SNDP}$, it is also feasible for path-relative SNDP [^1] Fix any $i\in[k]$ and $F\subseteq E$ with $|F|<r_{i}$. If GFG-F has an $s_{i}$-$t_{i}$ path, then $\delta_{G-F}(S)\neq\emptyset$ for every $s_{i}$-$t_{i}$ cut. So since $H$ is feasible for cut-relative $\mathsf{SNDP}$, we have $\delta_{H-F}(S)\neq\emptyset$ for every $s_{i}$-$t_{i}$ cut, which implies that HFH-F has an $s_{i}$-$t_{i}$ path. but the converse fails to hold; see Appendix A for an example.
@@
We arrived at $\mathsf{CR}\text{-}{f}\text{-}\mathsf{NDP}$, which can be seen as a fault-tolerant cut-covering problem, as a technically natural extension of $k\text{-}\mathsf{EFTS}$. We show below that $\mathsf{CR}\text{-}{f}\text{-}\mathsf{NDP}$ offers a surprising amount of modeling power, and, in particular, allows one to capture stronger forms of relative fault-tolerance compared to $k\text{-}\mathsf{EFTS}$. In $k\text{-}\mathsf{EFTS}$, there is a sharp relative-fault-tolerance threshold at $k$: if $H$ is feasible, then GFG-F and HFH-F have the same components whenever $|F|<k$, but there are no guarantees for larger fault-sets. With $\mathsf{CR}\text{-}{f}\text{-}\mathsf{NDP}$, one can capture a weaker guarantee also for other fault-sets, which allows for a more graceful degradation of relative fault-tolerance as $|F|$ increases.
@@
One way of modeling graceful relative-fault-tolerance degradation is as follows: given a non-increasing function $\tau:\mathbb{Z}_{+}\mapsto\mathbb{Z}_{+}$, we seek a min-cost subgraph $H$ such that for every fault-set $F$, GFG-F and HFH-F have exactly the same components with at most $\tau(|F|)$ nodes. If $G$ models a communication network where connected nodes can communicate with each other, then this yields the desirable guarantee that, post-faults, if a node can communicate with more than $\tau(|F|)$ nodes in $G$ then the same holds for $H$; if this number is at most $\tau(|F|)$, then post-faults, it can communicate with the same nodes in $G$ and $H$. Note that if $\tau(k-1)\geq n=|V|$, then $H$ is feasible for $k\text{-}\mathsf{EFTS}$; by adding more “levels” to $\tau$, we can obtain fault-tolerance guarantees for larger fault-sets.
@@
Generalizing Example 1.2 substantially, suppose along with $\tau$, we have a monotone function , i.e., if $T\subseteq S$. We now seek a subgraph $H$ such that for every fault-set $F$ and $S\subseteq V$ with , we have that $S$ is (the node-set of) a component of GFG-F iff it is a component of HFH-F. Similar to Example 1.2, we can model this as $\mathsf{CR}\text{-}{f}\text{-}\mathsf{NDP}$ by defining $f(S):=\min\,\{\ell:\tau(\ell)<$, for $S\neq\emptyset$ (Theorem 3.3).
This setup yields fault-tolerance degradation depending on monotone properties of components other than just their size, which creates a rich space of problems. For example, suppose is the weak-diameter of $S$. Setting $\tau(\ell)=n$ if $\ell<k$, and some $t<n$ for larger $\ell$, we seek a $k\text{-}\mathsf{EFTS}$ solution $H$ such that for $F\subseteq E$ with $|F|\geq k$, GFG-F and HFH-F have the same components of weak-diameter at most $t$.
@@ 1.1 Our contributions
We introduce a little notation to describe this. Let $\mathsf{CR}\text{-}{(h,D)}\text{-}\mathsf{NDP}$ denote $\mathsf{CR}\text{-}{f}\text{-}\mathsf{NDP}$ on graph $D$ with base cut-requirement function $h$. For $S\subseteq V$, define its deficiency to be $f(S)-|\delta_{G}(S)|$; we say that $S$ is a small cut if its deficiency is positive (i.e., $|\delta_{G}(S)|<f(S)$). We use G[S]G[{S}] to denote the subgraph induced by $S$. Let $\overline{S}$ denote VSV-S.
@@
Iterative rounding and relaxation derives its power from the fact that an extreme-point solution to an LP-relaxation of the problem can be defined using a structured family of tight constraints, such as, often a laminar family of cut constraints in the case of network-design problems. As noted earlier, this key property does not hold for cut-relative $\mathsf{SNDP}$ (and path-relative $\mathsf{SNDP}$). Recently, various works have considered network-design problems that give rise to cut constraints that are less structured and not quite uncrossable; see [2, 3, 6] and the references therein. Our work can be seen as furthering this research direction.
@@ 2 Preliminaries and notation
Recall that we are given an undirected graph $G=(V,E)$ with nonnegative edge costs $\{c_{e}\}_{e\in E}$. For a subset $S\subseteq V$ and subset $Z\subseteq E$, which we will interchangeably view as the subgraph $(V,Z)$ of $G$, we use $\delta_{Z}(S)$ to denote $\delta_{G}(S)\cap Z$, i.e., the edges of $Z$ on the boundary of $S$. Recall that $\overline{S}$ denotes VSV-S, and G[S]G[{S}] denotes the subgraph induced by $S$. A cut-requirement function on $G$ is a function $f:2^{V}\mapsto\mathbb{Z}$. (We allow $f(S)$ to be negative chiefly for notational simplicity; this does not affect anything, as we will only examine sets for which $f(S)\geq 0$.) Recall the following network-design problems.
@@ 4. $\bullet$
Let $f:2^{V}\mapsto\mathbb{R}$. We say that $f$ is weakly supermodular if for any two node-sets $A,B\subseteq V$, we have $f(A)+f(B)\leq\max\bigl{\{}f(A\cap B)+f(A\cup B),\,f(A-B)+f(B-A)\bigr{\}}$. We say that $f$ is symmetric if f(S)=f(VS)f(S)=f(V-S) for all $S\subseteq V$, and is normalized if $f(\emptyset)=f(V)=0$.
@@
    Let $X\subseteq V$, and $\mathcal{C}$ be a collection of subsets of VXV-X closed under taking set intersections, unions, differences, and complements with respect to VXV-X, i.e., if $S\in\mathcal{C}$ then $V-X-S\in\mathcal{C}$, and $S,T\in\mathcal{C}$ implies that $\{S\cap T,S\cup T,S-T,T-S\}\subseteq\mathcal{C}$.
@@ 1. $\bullet$
    $X=A$,  Y=VBY=V-B. We have $X\cap Y=A-B$, and $X\cup Y=V-(B-A)$, so $f^{\mathsf{sym}}(X\cup Y)=f^{\mathsf{sym}}(B-A)$. Also, $X-Y=A\cap B$, and $Y-X=V-(A\cup B)$, so $f^{\mathsf{sym}}(Y-X)=f^{\mathsf{sym}}(A\cup B)$. So (1) holds.
@@ 2. $\bullet$
    X=VAX=V-A,  $Y=B$. This is completely symmetric to the above case. We have $f^{\mathsf{sym}}(X\cap Y)=f^{\mathsf{sym}}(B-A)$ and $f^{\mathsf{sym}}(X\cup Y)=f^{\mathsf{sym}}(A-B)$. Also, $f^{\mathsf{sym}}(X-Y)=f^{\mathsf{sym}}(A\cup B)$ and $f^{\mathsf{sym}}(Y-X)=f^{\mathsf{sym}}(A\cap B)$.
@@ 3. $\bullet$
    X=VAX=V-A,  Y=VBY=V-B. We have $X\cap Y=V-(A\cup B)$, so $f^{\mathsf{sym}}(X\cap Y)=f^{\mathsf{sym}}(A\cup B)$, and $X\cup Y=V-(A\cap B)$, so $f^{\mathsf{sym}}(X\cup Y)=f^{\mathsf{sym}}(A\cap B)$. Also, XY=BAX-Y=B-A and YX=ABY-X=A-B.
@@
Now we show that $g_{2}$ is weakly supermodular. Consider $A,B\subseteq X$. If any of the sets $A-B,B-A,A\cap B,X-(A\cup B)$ are empty, then $g_{2}(A)+g_{2}(B)=g_{2}(A\cap B)+g_{2}(A\cup B)$ or $g_{2}(A)+g_{2}(B)=g_{2}(A-B)+g_{2}(B-A)$. The former holds when ABA-B or BAB-A is empty; the latter holds when $A\cap B$ or $X-(A\cup B)$ is empty, where we also utilize the fact that $g_{2}$ is symmetric. So assume otherwise. Then, for every $S\in\{A,B,A\cap B,A\cup B,A-B,B-A\}$, we have $g_{2}(S)=\max_{T\in\mathcal{C}}f(S\cup T)$.
@@ 2.1 LP-relaxation for $\mathsf{CR}\text{-}{f}\text{-}\mathsf{NDP}$
Fact 7 (Ellipsoid method [14]).
Given a polytime separation oracle for $(\text{{CRLP}}_{f,G})$, one can find an extreme-point optimal solution to LPs $(\text{{CRLP}}_{f,G})$ and (P’) in polytime.
@@
We show that for the $\mathsf{CR}\text{-}\mathsf{SNDP}$ instance shown in Fig. 1, the extreme-point solution $\widehat{x}$ shown in the figure, cannot be expressed as the solution to any laminar family of tight cut constraints (2) (along with tight edge constraints (3)). The $\mathsf{SNDP}$ instance is $s$-$t$ 22-edge connectivity: that is, the base cut-requirement function is given by $f(S)=2$ if $S$ is an $s$-$t$ cut, and $f(S)=0$ otherwise.
@@ 3 Modeling power of $\mathsf{CR}\text{-}{f}\text{-}\mathsf{NDP}$
    $H$ is feasible for $\mathsf{CR}\text{-}{f^{\mathsf{grace}}}\text{-}\mathsf{NDP}$ iff for every fault-set $F\subseteq E$, and every $S\subseteq V$ with $|S|\leq\tau(|F|)$, $S$ is (the node-set of) a component of GFG-F iff it is a component of HFH-F.
@@
To prove the feasibility characterization in part (b), we note that the function $f^{\mathsf{grace}}$ is constructed so that for any $S\subseteq V$ and $F\subseteq E$, we have $|S|\leq\tau(|F|)$ iff $f^{\mathsf{grace}}(S)>|F|$, so that $|\delta_{H}(S)|\geq f^{\mathsf{grace}}(S)$ implies that $S$ is not a component of HFH-F. Suppose $H$ is such that $|\delta_{H}(S)|\geq\min\bigl{\{}f^{\mathsf{grace}}(S),|\delta_{G}(S)|\bigr{\}}$ for all $S\subseteq V$. Consider any fault-set $F\subseteq E$ and any $S\subseteq V$ such that $|S|\leq\tau(|F|)$. By the above, if $S$ is a component of HFH-F, then $|\delta_{H}(S)|<f^{\mathsf{grace}}(S)$, and so $S$ must also be a component of GFG-F. But this also implies that if $S$ is a component of GFG-F, then it is a component of HFH-F; otherwise, there is some $A\subsetneq S$ that is a component of HFH-F with $|A|<|S|\leq\tau(|F|)$, and so $A$ must be a component of GFG-F, which is not the case.
Conversely, suppose that for every $F\subseteq E$, and $S\subseteq V$ with $|S|\leq\tau(|F|)$, we have that $S$ is a component of GFG-F iff it is a component of HFH-F. If $|\delta_{H}(S)|<\min\bigl{\{}f^{\mathsf{grace}}(S),|\delta_{G}(S)|\bigr{\}}$ for some $S\subseteq V$, then take $F=\delta_{H}(S)$. Since $f^{\mathsf{grace}}(S)>|F|$, we have $|S|\leq\tau(|F|)$. But $S$ is a component of HFH-F and not GFG-F, which yields a contradiction.
@@
    $H$ is feasible for $\mathsf{CR}\text{-}{f}\text{-}\mathsf{NDP}$ iff for every fault-set $F\subseteq E$, and every $S\subseteq V$ with , $S$ is (the node-set of) a component of GFG-F iff it is a component of HFH-F.
@@
For part (b), similar to before, we have that iff $f(S)>|F|$. Suppose $H$ is feasible for $\mathsf{CR}\text{-}{f}\text{-}\mathsf{NDP}$. Consider any fault-set $F\subseteq E$ and any $S\subseteq V$ be such that . By the above, $f(S)>|F|$, so if $S$ is a component of HFH-F, then it must also be a component of GFG-F. Again, this also implies that if $S$ is a component of GFG-F, then it is a component of HFH-F: otherwise, there is some $A\subsetneq S$ that is a component of HFH-F, and , which means that $A$ is a component of GFG-F.
Conversely, suppose that for every $F\subseteq E$, and $S\subseteq V$ with , we have that $S$ is a component of GFG-F iff it is a component of HFH-F. If $|\delta_{H}(S)|<\min\bigl{\{}f(S),|\delta_{G}(S)|\bigr{\}}$ for some $S\subseteq V$, then take $F=\delta_{H}(S)$. Since $f(S)>|F|$, we have . But $S$ is a component of HFH-F and not GFG-F, which yields a contradiction.